OpenStudy (anonymous):

is (-1,-1) (1,3) equal 3--1 over 1--1 = 4/2 = 2?

OpenStudy (anonymous):

\[\left[\begin{matrix}3 & +1 \\ 1& +1\end{matrix}\right] = \left(\begin{matrix}4 \\ 2\end{matrix}\right)\]

OpenStudy (anonymous):

which equals 2

OpenStudy (anonymous):

you mean slope right?

OpenStudy (anonymous):

he got me confused here!!

OpenStudy (anonymous):

its just those y2-y1 over x2-x1 equations

OpenStudy (anonymous):

(-1,-1) (1,3) represents 2 points on a graph right?

OpenStudy (anonymous):

wait.. let me rephrase.. forget everything i said.. Write the equation of the line through the points (-1,-1), (1,3)

OpenStudy (anonymous):

ok! now what you worked out is the slope the line = 2 now you use the standard equation y-y1 = m(x-x1) where m = slope and (x1,y1) is a point on the line we can use any of the two points in your question lets use (1,3) x1 = 1 and y1 = 3 we get: y - 3 = 2(x-1) y - 3 = 2x - 2 y = 2x + 1 - theres the equation of your line